# Composed Multi-modal Retrieval: A Survey of Approaches and Applications

### v1

@@ 6.5 Composed Person Retrieval
Composed Person Retrieval (CPR) represents an innovative approach to identifying specific individuals by leveraging both visual and textual information. Traditional methods, such as Image-based Person Retrieval (IPR) and Text-based Person Retrieval (TPR), often fall short in effectively utilizing both types of data, leading to a loss in accuracy. CPR aims to address this limitation by simultaneously employing image and text queries to enhance the retrieval process. This dual-modality approach not only increases the descriptive power of the query but also refines the relevance of search results, providing more accurate identification of target individuals. CPR is particularly useful in various applications, including social services and public security, where precise person identification is crucial.
The introduction of CPR has significant practical implications, offering substantial benefits for both business and social sectors. By improving the efficiency and accuracy of person retrieval, CPR enhances decision-making processes and user experiences through tailored solutions. For instance, in fields ranging from law enforcement to personalized marketing, CPR provides exceptional data processing capabilities and robust support for related research activities.
@@ 6.5.1 Benchmark Datasets
ITCPR [100] is a meticulously curated dataset designed for evaluating the performance of Zero-shot Composed Person Retrieval (ZS-CPR). It consists of finely annotated image-text pairs specifically created to address the challenges posed by CPR without requiring costly manual annotations. ITCPR leverages existing image-text datasets relevant to TPR tasks, ensuring a broad coverage of diverse scenarios. The dataset includes over 21,000 queries, each paired with corresponding images that depict the same identity under different outfits or conditions, thereby forming complete triplets (Iq, Tq, It). In ITCPR, reference images are combined with relative captions to form multimodal queries. These queries are designed to capture subtle variations in appearance, such as clothing changes or accessory additions, which are essential for distinguishing between individuals. Each triplet in the dataset comprises a reference image (Iq), a relative caption (Tq), and several target images (It) depicting the same individual in different contexts.
@@ 6.6 Composed Sketch-based Image Retrieval
The objective of Composed Sketch-Text Image Retrieval [101] is to combine sketch and textual information to enhance the accuracy and relevance of image retrieval. Specifically, it aims to leverage sketches to capture the shape and structure of objects, while textual descriptions supplement visual details such as color, material, and texture. This approach improves retrieval by integrating both coarse and fine-grained visual information. For instance, users may express their needs by sketching or describing an image, even in the absence of a specific picture.
The application value and practical use cases are diverse, including but not limited to the following areas: (1) In the design and creative industries, designers can use sketches to convey concepts and pair them with textual descriptions to aid in searching for inspiration or reference material within large databases, particularly in fields like fashion design and interior decoration. (2) In e-commerce and online shopping, consumers can provide hand-drawn sketches or textual descriptions of a product, allowing the system to recommend similar product images, thereby enhancing the shopping experience. (3) In healthcare, doctors can use this technology to identify rare diseases or medical devices by sketching the affected area or device shape and supplementing it with descriptive text, allowing for rapid location of relevant information to support diagnostic decisions. (4) In law enforcement, police can search for suspects based on eyewitness descriptions and suspect sketches, thereby improving investigative efficiency.
@@ 6.6.1 Benchmark Datasets
QMUL-ShoeV2 and QMUL-ChairV2 [102]: Both datasets emphasize detailed associations between sketch queries and corresponding images, with QMUL-ShoeV2 comprising 2000 sketches and 6730 photos, while QMUL-ChairV2 includes 400 sketches and 1800 photos.
@@ 6.7 Interactive/Conversational Retrieval
Interactive/Conversational Retrieval (ICR) represents an advanced approach to image retrieval that leverages natural language interactions between users and systems to refine search outcomes progressively [84, 109, 110, 111, 112, 113, 114, 115, 116, 117]. Unlike traditional methods relying solely on images or predefined textual attributes, ICR integrates user feedback through conversational interfaces, enhancing the accuracy and relevance of search results. This method enables users to provide iterative feedback in natural language, refining queries dynamically until they locate the desired image or item. The primary objective of ICR is to facilitate more intuitive, precise, and personalized searches by incorporating both visual and semantic information effectively.
ICR has significant applications across various domains, including e-commerce, fashion, and social media, where users often seek specific items with complex characteristics such as style, color, and material. By enabling multi-turn interactions, ICR improves user experience and satisfaction by allowing for more detailed and nuanced queries. For instance, in fashion retail, customers can start with a reference image of a clothing item and iteratively modify their query (”long-sleeve version”, ”in blue color”) to find exactly what they are looking for. This not only enhances the shopping experience but also supports retailers in providing accurate product recommendations based on user preferences, thereby boosting customer engagement and sales.
@@ 6.7.1 Benchmark Datasets
Several benchmark datasets have been developed to evaluate and advance the capabilities of Interactive/Conversational Retrieval systems. These include:
@@
Multi-turn Shoes: The initial Shoes dataset [118] encompasses a collection of images featuring women’s footwear, spanning 10 distinct categories, which were sourced online and automatically annotated with various attributes. To enhance the applicability of these images for single-turn, feedback-oriented image retrieval tasks, [118] introduced supplementary natural language descriptions, creating about 10k training pairs and 4.6k test queries. Within the scope of this work, Pal et al. [113] extended this dataset to accommodate multi-turn interactions, linking multiple single-turn interactions by matching the target image from one session to the query image of another. This extension facilitates research into multi-turn retrieval scenarios by concatenating several single-turn transactions and maintaining consistency with Multi-turn FashionIQ, thus offering valuable insights into memory retention and feedback handling across multiple turns.
@@
Interactive Retrieval [110] is meticulously designed to support multi-turn interactive retrieval tasks involving complex scenes. Originating from the Visual Genome [krishna2017visual], this dataset includes detailed captions for various regions within each image, making it ideal for scenarios where users refine queries based on feedback. Comprising 8,960 training samples, approximately 10% of the entire dataset, it evaluates model performance under limited labeled data conditions. For rigorous assessment, a subset featuring over 18 nouns per caption, totaling 7,000 samples, emphasizes challenging cases with potential confounders and spurious correlations. A user simulator mimics real human interactions by providing image-click and text-click feedback on candidate images. Validated through 2,500 user sessions conducted by six annotators, the simulator accurately predicts actual user clicks in 85% and 71.4% of sessions, respectively.

### v2

@@ 6.5 Composed Person Retrieval
Composed Person Retrieval (CPR) represents an innovative approach to identifying specific individuals by leveraging both visual and textual information. Traditional methods, such as Image-based Person Retrieval (IPR) [255] and Text-based Person Retrieval (TPR) [256, 257], often fall short in effectively utilizing both types of data, leading to a loss in accuracy. CPR aims to address this limitation by simultaneously employing image and text queries to enhance the retrieval process. This dual-modality approach not only increases the descriptive power of the query but also refines the relevance of search results, providing more accurate identification of target individuals. CPR is particularly useful in social services and public security, where precise person identification is crucial.
@@ 6.5.1 Benchmark Datasets
SynCPR [111] contains 1.15 million high-quality triplets, making it one of the largest synthetic datasets for CPR tasks. It includes diverse scenarios, broad age coverage, and comprehensive ethnic representation, ensuring high quality, realism, and diversity of person images.
ITCPR [111] contains 2,225 annotated triplets, comprising 2,202 unique query combinations, from 1,199 identities. The gallery consists of 20,510 person images, among which 2,225 correspond directly to queries. The dataset is carefully reviewed to eliminate potential false-negative cases, ensuring reliable evaluation metrics. The textual annotations have an average sentence length of 9.54 words, with the longest sentence containing 32 words and the shortest containing 3 words. This dataset is exclusively designated for testing in the ZS-CPR task.
@@ 6.5.2 Results and Analysis
We report the evaluation results of Composed Person Retrieval on the ITCPR dataset, as shown in Table XII. All models are pre-trained on the large-scale SynCPR dataset and tested under the zero-shot setting. Among the compared methods, FAFA [111] achieves the best performance, reaching 46.54% in Recall@1 and 55.60% in mAP, surpassing both CaLa [60] and SPRC [63]. These results demonstrate the effectiveness of incorporating fine-grained fusion strategies and synthetic data in improving retrieval accuracy.
@@ 6.6 Composed Sketch-based Image Retrieval
Composed Sketch-Text Image Retrieval [112] aims to improve the accuracy and relevance of image retrieval by integrating sketch-based and textual inputs. This approach leverages sketches to capture object shapes and structures, while textual descriptions provide complementary details such as color, material, and texture. By combining coarse structural information with fine-grained attributes, it enables more expressive and flexible querying, especially useful when users lack a specific reference image.
This method offers broad application potential across various domains. In design and creative industries, designers can use sketches paired with text to search for inspiration or references in large databases, supporting fields such as fashion design and interior decoration. In e-commerce, consumers can input hand-drawn sketches or textual descriptions to find visually similar products, enhancing the online shopping experience. In law enforcement, investigators can use textual descriptions and suspect sketches to improve the efficiency of suspect identification and criminal investigations.
@@ 6.6.1 Benchmark Datasets
ShoeV2 and ChairV2 [154]: Both datasets emphasize associations between sketch queries and corresponding images, with ShoeV2 comprising 2000 sketches and 6730 photos, while ChairV2 includes 400 sketches and 1800 photos.
@@ 6.6.2 Results and Analysis
We report retrieval results on five widely used benchmark datasets (Table XIII), including ShoeV2, ChairV2, Sketchy, FS-COCO, and SketchyCOCO. Across all datasets, the recently proposed SCIR method [112] achieves the best performance, with Recall@10 reaching 81.4% on ChairV2 and 79.1% on ShoeV2. Notably, SCIR also leads on more complex datasets such as FS-COCO and SketchyCOCO, which require the model to jointly reason over sketch-based structural information and textual semantics. While earlier methods like Combiner and SEARLE incorporate both modalities, their fusion strategies are relatively shallow and offer limited cross-modal interaction, particularly in capturing fine-grained attributes. In contrast, SCIR introduces deeper alignment and richer interactions between sketch and text representations, resulting in substantial performance improvements, exceeding 10% gain in Recall@10 on several benchmarks. These results underscore the importance of well-designed multimodal fusion mechanisms for retrieving images that require both global contour understanding and fine-grained attribute reasoning.
@@ 6.7 Interactive/Conversational Retrieval
Interactive/Conversational Retrieval (ICR) represents an advanced approach to image retrieval that leverages natural language interactions between users and systems to refine search outcomes progressively [28, 113, 114, 115, 116, 117, 118, 119, 120, 121]. Unlike traditional methods relying solely on images or predefined textual attributes, ICR integrates user feedback through conversational interfaces, enhancing the accuracy and relevance of search results. This method enables users to provide iterative feedback in natural language, refining queries dynamically until they locate the desired image or item. The primary objective of ICR is to facilitate more intuitive, precise, and personalized searches by incorporating both visual and semantic information effectively. ICR has significant applications across various domains, including e-commerce, fashion, and social media.
@@ 6.7.1 Benchmark Datasets
Several benchmark datasets have been developed to evaluate and advance the capabilities of Interactive Conversational Retrieval systems. These include:
@@
Multi-turn Shoes [117]: The initial Shoes dataset [259] encompasses a collection of images featuring women’s footwear, spanning 10 distinct categories, which were sourced online and automatically annotated with various attributes. To enhance the applicability of these images for single-turn, feedback-oriented image retrieval tasks, work [259] introduced supplementary natural language descriptions, creating about 10k training pairs and 4.6k test queries. Within the scope of this work, Pal et al. [117] extended this dataset to accommodate multi-turn interactions, linking multiple single-turn interactions by matching the target image from one session to the query image of another. This extension facilitates research into multi-turn retrieval scenarios by concatenating several single-turn transactions and maintaining consistency with Multi-turn FashionIQ, thus offering valuable insights into memory retention and feedback handling across multiple turns.
@@
Interactive Retrieval [114] is meticulously designed to support multi-turn interactive retrieval tasks involving complex scenes. Originating from the Visual Genome [260], this dataset includes detailed captions for various regions within each image, making it ideal for scenarios where users refine queries based on feedback. Comprising 8,960 training samples, approximately 10% of the entire dataset, it evaluates model performance under limited labeled data conditions. For rigorous assessment, a subset featuring over 18 nouns per caption, totaling 7,000 samples, emphasizes challenging cases with potential confounders and spurious correlations. A user simulator mimics real human interactions by providing image-click and text-click feedback on candidate images. Validated through 2,500 user sessions conducted by six annotators, the simulator accurately predicts actual user clicks in 85% and 71.4% of sessions, respectively.
6.7.2 Results and Analysis
In existing multi-turn conversational retrieval studies, many methods provide only qualitative analyses, such as visualizations or case-based discussions, without reporting quantitative performance metrics. Therefore, we summarize a few representative approaches with available results, as shown in Table XIV. We report quantitative retrieval performance on the Multi-turn FashionIQ test set (Table XIV). In this task, FashionNTM [117] achieves the highest Recall@5 and Recall@8 across all subcategories (Dress, Shirt, Toptee), significantly outperforming earlier methods. For example, it reaches a Recall@5 of 48.3% in the Dress category, indicating a clear overall performance advantage. This remarkable performance is likely due to its novel architecture based on a Cascaded Memory Neural Turing Machine (CM-NTM), which effectively manages the conversational state and integrates feedback from all previous turns to model the evolving user intent. In terms of evaluation metrics, in addition to the widely used Recall@K, we adopt MRR (Mean Reciprocal Rank) to better capture retrieval efficiency in interactive settings. MRR calculates the reciprocal of the rank at which the target image first appears and then averages this value over all queries. A higher MRR indicates that the system tends to retrieve relevant results earlier in the ranked list, which is particularly important in iterative feedback scenarios. Additionally, IRR [118] and CAFA [115] also demonstrate competitive performance, with Recall@5 exceeding 25% and MRR reaching 20.6% and 26.5%, respectively. In contrast, earlier methods such as TIRG and ComposeAE show limited capacity in handling multi-turn natural language feedback.
7 FUTURE DIRECTIONS
Composed multi-modal retrieval has emerged as a rapidly evolving research field that, while achieving significant progress, continues to face numerous challenges and opportunities. Current CMR techniques, although demonstrating excellent performance under supervised learning paradigms, still exhibit notable limitations in zero-shot generalization, complex semantic understanding, and real-time performance. Concurrently, the rapid advancement of generative artificial intelligence, LLMs, and edge computing technologies provides novel technical approaches and solutions for CMR research. Furthermore, practical application scenarios demand higher requirements for personalization, explainability, and privacy protection, driving the field toward more practical and industrialized development. Based on current technological trends and practical application needs, we believe future research should focus on the following directions:
Generative Data Construction and Quality Control. (1) Controllable Synthetic Data Generation: The primary challenge facing current CMR methods is the insufficient quality and diversity of generated data. Future research needs to develop more refined generation control mechanisms, utilizing diffusion models and generative adversarial networks to achieve precise control over modification types, semantic intensity, and visual attributes. (2) Multi-level Data Quality Assessment: Constructing automated quality assessment frameworks that encompass multiple dimensions, including semantic consistency, visual rationality, and textual accuracy. By combining pre-trained multi-modal discriminators with human feedback reinforcement learning techniques, real-time quality monitoring and filtering of synthetic data can be achieved, thereby substantially improving the overall quality of training data. (3) Hard Negative Data Augmentation: Designing targeted hard sample generation strategies by constructing negative samples that are semantically similar but exhibit significant detail differences, thereby enhancing model discriminative capabilities. This approach can effectively alleviate the insufficient distinction between positive and negative samples in existing datasets, strengthening model robustness in complex retrieval scenarios.
Complex Query Understanding Reasoning. (1) Temporal and Causal Reasoning: Extending CMR capabilities to handle temporal relationships and causal logic, which is particularly important for video retrieval and dynamic scene understanding. (2) Common-sense Knowledge Integration: Integrating large-scale knowledge graphs and common-sense reasoning capabilities into CMR systems, enabling models to understand implicit semantic associations. (3) Multi-turn Interaction Optimization: Developing conversational retrieval systems with memory capabilities capable of understanding cross-turn semantic dependencies and user intent evolution. Through maintaining dialogue states, understanding elliptical expressions, and handling coreference resolution, more natural and fluent interactive experiences can be provided.
Explainability and System Trustworthiness. (1) Multi-granularity Explanation Generation: Developing CMR systems capable of providing explanations at different abstraction levels. From fine-grained feature activation visualization to medium-grained semantic region annotation and high-level decision-logic explanation, corresponding explanatory information can be provided for users with different needs. This is significant for enhancing user trust and system transparency. (2) Uncertainty Quantification: Introducing Bayesian deep learning and other techniques to quantitatively assess the reliability of retrieval results. When models have an ambiguous understanding of certain queries, they should actively seek clarification from users or provide multiple candidate explanations for user selection. (3) Adversarial Robustness: Enhancing CMR system robustness against input perturbations and adversarial attacks. In particular, in security-sensitive application scenarios, the focus is on ensuring that systems maintain stable performance when faced with malicious input.
Efficiency Optimization and Real-time Deployment. (1) Model Compression: Developing lightweight, specialized CMR models. Through knowledge distillation, neural network pruning, and quantization techniques, model complexity can be significantly reduced while maintaining retrieval accuracy. Particularly for mobile and edge devices, adaptive model architectures are needed that can dynamically adjust computational complexity based on the device. (2) Hierarchical Retrieval Architecture: Designing multi-level retrieval systems that combine coarse-grained pre-filtering with fine-grained precise matching to substantially improve retrieval efficiency for large-scale databases. While ensuring retrieval quality, millisecond-level response speeds can be achieved to meet real-time application requirements.
Cross-modal Extension and Emerging Applications. (1) Immersive Experience Applications: Expanding from current image-text combinations to richer modal combinations including audio, video, 3D models, and sensor data. For example, users can quickly retrieve and customize 3D objects, scene layouts, and other content in virtual environments through natural language descriptions and gesture interactions. (2) Scientific Research Support: Applying CMR technology to scientific data analysis, such as “finding images similar to reference cases but with clearer lesions” in medical image diagnosis, or “searching for materials with similar crystal structures but different doping elements” in materials science. These applications require highly specialized domain knowledge integration.
Standardized Evaluation and Fairness. (1) Multi-dimensional Evaluation Framework: Establishing more comprehensive evaluation systems that consider not only retrieval accuracy but also response time, user satisfaction, diversity, and other dimensions. Designing specialized evaluation protocols for different application scenarios to ensure consistency between evaluation results and actual application effectiveness. (2) Bias Detection and Mitigation: Establishing systematic bias detection mechanisms to identify and mitigate algorithmic biases across dimensions such as gender, race, and age. Through fairness constraints and adversarial debiasing techniques, ensure that retrieval systems provide fair services for all user groups.